# Retrieval Enhanced Zero-Shot Video Captioning

### v1

@@ 3.3. Inference
Optimization occurs throughout the sentence generation, enhancing the correspondence between the generated sentence and video through the iterative application. More specifically, the optimization is conducted at each step in the auto-regression process. When generating the dot token or reaching the maximum sentence generation length, we stop the process and initiate the generation of a new sentence while retaining the optimized soft prompt.
@@
After all iterations are complete, we obtain $M$ generated captions. We calculate the similarity between these $M$ generated captions and the video as described in Eq. (6). The caption with the highest similarity is selected as the final caption.
4. Experiments
4.1. Datasets and Metrics
@@
MSR-VTT (Xu et al., 2016) is a large-scale video captioning dataset collected from a commercial video website. It contains 10,000 video clips covering 20 categories. Each clip has 20 English annotations. In the standard split, it includes 6,513 training videos, 497 validation videos, and 2,990 test videos.
@@
MSVD (Chen and Dolan, 2011) is a widely-use video captioning dataset, which is collected from Youtube. It consists of 1,970 video clips. Each clip has roughly 40 ground-truth captions. In the standard split, it includes 1,200 training videos, 100 validation videos, and 670 test videos.
@@
VATEX (Wang et al., 2019) is a large-scale dataset that contains 34,991 videos with 10 English annotations. The standard split includes 25,910 training videos, 3,000 validation videos, and 6,000 test videos.
@@
Metrics. In our experiments, we adopt four commonly used metrics to evaluate caption quality, including BLEU@4 (B@4) (Papineni et al., 2002), METEOR (M) (Banerjee and Lavie, 2005), ROUGE-L (R) (Lin and Och, 2004), and CIDEr (C) (Vedantam et al., 2015). Among these metrics, CIDEr is the main metric.
4.2. Implementation Details
@@
We use the pre-trained XCLIP [^1] https://huggingface.co/microsoft/xclip-base-patch16-kinetics-600-16-frames as our video retriever. We uniformly sample 16 frames per video for video retrieval. The visual feature extraction and text-image similarity calculation are performed using the pre-trained CLIP ViT-L/14222https://github.com/openai/CLIP model. We first sample three frames every second for the given video. To avoid using duplicate frames, we further use the CLIP-based strategy to subsample the keyframes. We start with the first frame as an anchor. The next frame is included in the set $\mathcal{F}$ if its dot product with the anchor, calculated by the CLIP image encoder, exceeds a specific threshold $\lambda_{CLIP}$. The newly sampled frame becomes the new anchor and the sampling process repeats in the same manner. The $\lambda_{CLIP}$ is set to 0.9 in all experiments. We use the pre-trained GPT-2 medium [^3] https://huggingface.co/openai-community/gpt2-medium as our caption generator, and the pre-trained AngLE [^4] https://github.com/SeanLee97/AnglE to calculate the similarity between the generated caption and retrieval sentences. The aforementioned pre-trained models are publicly and readily available.
@@
Empirically, $\lambda_{l}$, $\lambda_{v}$, $\lambda_{s}$ and $\lambda_{w}$ are set to 1.6, 1.0, 0.8, and 0.3 respectively. In our method, we set the number of learnable tokens to 5. We set the number of retrieval sentences to 15 and the frequency size to 5. We adopt the AdamW (Loshchilov and Hutter, 2019) optimizer with a learning rate of 1e-4 and a weight decay weight of 0.3 to optimize our model. Throughout the experiments, we employ 16 generation iterations per video. To prevent sentences from being overly long and repetitive, we limit the number of tokens generated per sentence to 15. We implement our framework with PyTorch and all experiments are conducted on 4 $\times$ RTX-A6000 GPU.
@@
Table 1. Results on MSRVTT across various models and supervision levels.
@@
VNS-GRU (Chen et al., 2020) Supervised 45.3 29.9 63.4 53.0
SemSynAN (Perez-Martin et al., 2021) 46.4 30.4 64.7 51.9
HMN (Ye et al., 2022) 43.5 29.0 62.7 51.5
SGN (Ryu et al., 2021) 40.8 28.3 60.8 49.5
SAAT (Zheng et al., 2020) 40.5 28.2 60.9 49.1
ZeroCap (Tewel et al., 2022b) Zero-shot 2.9 16.3 35.4 9.6
MAGIC (Su et al., 2022) 5.5 13.3 35.4 7.4
EPT(Tewel et al., 2022a) 3.0 14.6 27.7 11.3
Video-LLaMA(Zhang et al., 2023) 4.9 16.8 25.3 2.3
ZS-CapCLIP(Radford et al., 2021) 4.0 15.0 31.0 5.0
MultiCapCLIP(Yang et al., 2023) 13.3 19.5 43.3 15.5
DeCap-BookCorpus (Li et al., 2023) 6.0 12.7 - 12.3
DeCap-CC3M (Li et al., 2023) 6.2 14.9 - 15.0
DeCap-COCO (Li et al., 2023) 14.7 20.4 - 18.6
Ours 3.5 17.3 28.7 22.1
Table 2. Results on MSVD across various models and supervision levels.
Methods Setting Metrics
B@4 M R C
Results on MSVD test set
VNS-GRU (Chen et al., 2020) Supervised 66.5 42.1 79.7 121.5
SemSynAN (Perez-Martin et al., 2021) 64.4 41.9 79.5 111.5
HMN (Ye et al., 2022) 59.2 37.7 75.1 104.0
ZeroCap (Tewel et al., 2022b) Zero-shot 2.3 12.9 35.4 9.6
MAGIC (Su et al., 2022) 6.6 16.1 40.1 14.0
EPT(Tewel et al., 2022a) 3.0 17.8 31.4 17.4
Video-LLaMA(Zhang et al., 2023) 4.0 15.1 21.6 1.3
Ours 5.0 24.0 36.9 38.1
Table 3. Results on VATEX across various models and supervision levels.
@@
VaTeX (Wang et al., 2019) Supervised 28.4 21.7 - 45.1
HMN (Ye et al., 2022) 33.2 22.7 49.3 51.6
DeCap-BookCorpus (Li et al., 2023) Zero-shot 4.1 9.9 - 11.8
DeCap-CC3M (Li et al., 2023) 7.3 12.6 - 18.4
DeCap-COCO (Li et al., 2023) 13.1 15.3 - 18.7
Video-LLaMA(Zhang et al., 2023) 4.3 16.3 21.8 3.8
Ours 2.8 14.1 23.5 23.9
4.3. Performance Comparison
@@
We compare our method with several state-of-the-art models under different settings on three public benchmark datasets. The results are shown in Table 1, 2 and 3. We note that the performance of supervised algorithms is significantly higher than zero-shot methods, but they require expensive annotation costs. Compared to the other zero-shot methods, our method achieves the best performance in terms of the main metric CIDEr on all the three datasets. As shown in Table 2, our method achieves a CIDEr score of 38.1 on the MSVD benchmark, leading by about an absolute 20.7 over the previous best result. On the more challenging benchmark MSR-VTT, our method also boosts the CIDEr score by about 18.8%. On the VATEX dataset, our method achieves a 23.9 CIDEr Score, gaining 18.1% improvement in terms of the main metric CIDEr on VATEX.
4.4. Ablation Study
Table 4. Results under different numbers of retrieved sentences.
Dataset # $K$ B@4 M R C
MSR-VTT 1 5 3.4 16.9 28.1 20.9
2 10 3.3 17.0 28.2 21.7
3 15 3.5 17.3 28.7 22.1
4 20 3.0 17.1 28.4 21.5
MSVD 5 5 5.2 23.4 36.3 36.0
6 10 4.7 23.8 36.2 37.3
7 15 5.0 24.0 36.9 38.1
8 20 4.9 23.9 36.1 37.5
VATEX 9 5 2.7 13.4 22.5 23.1
10 10 3.0 13.8 22.9 23.5
11 15 2.8 14.1 23.5 23.9
12 20 2.7 13.9 23.4 23.3
@@ 4.4. Ablation Study
Effect of the number of retrieved sentences. We study the effect of using different numbers of retrieved sentences. Table 4 shows the difference in video captioning performance as $K$ varies from 5 to 20. We find that retrieving 15 sentences offers better CIDEr scores across three datasets compared to other numbers. It can be concluded that too many or too few retrieval sentences can degrade the performance. This is because too many retrieved sentences may introduce too much noise while too few retrieved sentences cannot provide sufficient information.
@@
Table 5. Effect of retrieval corpora. The experiments are conducted on test split of MSVD, MSR-VTT and VATEX datasets.
@@
MSR-VTT 1 TrainSet 2.5 16.6 27.0 18.8
2 TestSet 3.5 17.3 28.7 22.1
MSVD 3 TrainSet 4.3 21.9 33.7 30.5
4 TestSet 5.0 24.0 36.9 38.1
VATEX 5 TrainSet 2.6 13.0 21.9 21.5
6 TestSet 2.8 14.1 23.5 23.9
Effect of different retrieval corpora. In Table 5, we compare the performance of our method under different retrieval corpora. On the MSR-VTT benchmark, the CIDEr score increases around $17.5\%$ by leveraging the TestSet as the retrieval corpus instead of the TrainSet (Row 1 v.s. Row 2). On MSVD and VATEX datasets, employing TestSet improves around $18.0\%$ (Row 3 v.s. Row 4) and $11.2\%$ (Row 5 v.s. Row 6) compared to TrainSet. This demonstrates that different knowledge corpora can significantly impact performance. In this paper, unless otherwise specified, we use TestSet as our default text corpus.
Effect of video retriever. In order to study the impact of video retrievers on our method, we use different configurations of XCLIP as our retriever. The results are present in Table 6. We observe that when XCLIP utilizes a ViT (Dosovitskiy et al., 2021) with a size of B/16 as the backbone, it achieves an approximate improvement of 1.4% in the CIDEr score compared to using B/32 (Row 2 v.s. Row 1). The Kinetics-600 (Li et al., 2020) dataset is the extended version of Kinetics-400 (Kay et al., 2017), containing more training data. Therefore, XCLIP pre-trained on Kinetics-600 brings a further improvement of 3.2% compared to XCLIP pre-trained on Kinetics-400 (Row 3 v.s. Row 2). Row 4 achieves the best performance on all metrics. It is trained with 16 frames per video and outperforms Row 3 using 8 frames. Therefore, in the other experiments conducted in this paper, we use the XCLIP in Row 4 as our default retriever.
Table 6. Performance with different configuration XCLIP. The experiments are conducted on the MSVD dataset.
@@
1 VIT-B/32 8 Kinetics-400 4.5 23.2 36.2 35.9
2 VIT-B/16 8 Kinetics-400 4.5 23.4 36.3 36.4
3 VIT-B/16 8 Kinetics-600 4.7 23.8 36.7 37.6
4 VIT-B/16 16 Kinetics-600 5.0 24.0 36.9 38.1
@@
Figure 3. Performance (CIDEr and BLEU@4 score) comparison under different numbers of soft tokens on the MSVD dataset.
Table 7. Performance with different granularity of retrieval loss. Experiments are conducted on test split.
Dataset # Prefix $\mathcal{L}_{retrieval}^{S}$ $\mathcal{L}_{retrieval}^{W}$ B@4 M R C
MSR-VTT 1 2.4 14.3 26.0 10.7
2 2.4 16.9 27.0 21.4
3 3.1 16.6 27.9 17.8
4 3.5 17.3 28.7 22.1
5 1.4 11.2 20.5 7.1
MSVD 6 2.8 17.3 30.0 16.8
7 4.3 22.9 34.7 37.2
8 4.8 22.5 34.6 30.6
9 5.0 24.0 36.9 38.1
10 2.0 13.8 23.4 10.4
VATEX 11 1.4 9.8 19.4 8.9
12 1.7 13.5 21.4 23.0
13 2.7 12.7 22.4 18.9
14 2.8 14.1 23.5 23.9
15 1.3 7.6 13.3 7.3
Effect of our proposed retrieval loss. To fully analyze the effect of our proposed two granularities of retrieval loss, we conduct experiments on MSR-VTT, MSVD and VATEX. As shown in Table 7, we observe that each loss assists the baseline to get better performance across the three datasets. Taking the results on MSR-VTT as an example, Row 2 shows that using only sentence-level retrieval loss leads to a CIDEr score increase of $100\%$ compared to the baseline in Row 1. In contrast, employing word-level retrieval loss results in a CIDEr enhancement of $66.3\%$ and a BELU@4 score improvement of $29.1\%$ (Row 3 v.s. Row 1). The best performance is achieved when both types of losses are used together (Row 4 v.s. Row 1). In Row 5, we also show the performance using the retrieved sentences as prefixes, which drops significantly. On the other datasets, our designed loss also achieves excellent performance. For MSVD, using sentence-level retrieval loss in Row 7 and word-level retrieval loss in Row 8 individually resulted in approximately 121.4% and 82.1% improvements in CIDEr score, respectively, compared to Row 6. When combining the two granularity losses, the CIDEr score increases by about 126.7% (Row 9 v.s.  Row 6). Row 10 shows that using the retrieved sentences as prefixes on MSVD also results in a significant performance decline. On the VATEX benchmark, sentence-level retrieval loss contributes more than word-level retrieval loss (Row 12 v.s.  Row 11, Row 13 v.s.  Row 11). In the comparison, the full model reaches the biggest improvement of $168.5\%$ in the term of CIDEr (Row14 v.s.  Row11). In Row 15, using the retrieved text as a prefix leads to the worst performance on VATEX.
Effect of the frequency size of the word-granularity loss. In Fig. 4, we conduct experiments on MSVD to explore the impact of the frequency size $L$ of word-granularity loss. From the figure, we observe that the CIDEr score first rises before reaching the moderate number (i.e., $L=5$), and then starts to decrease slightly. The main reason is that when $L$ is large, some noisy words are introduced. On the contrary, if $L$ is small, some essential words may be neglected. In Fig. 4(b), a similar trend is observed in the BLEU metric as well. Therefore, we select an appropriate number (i.e., $L=5$) as the frequency size.
@@
Figure 4. Performance (CIDEr and BLEU@4 score) comparison under different frequency sizes on the MSVD dataset.
Figure 5. Several video captioning qualitative examples using different methods.
Figure 6. Several image captioning qualitative examples using different methods.
Zero-shot image captioning. Our method can be adapted for zero-shot image captioning by swapping XCLIP with CLIP. In Table 8, we compare the performance among several zero-shot image captions methods on the MS-COCO (Chen et al., 2015) dataset. We achieve the second-best results on all metrics, only falling behind MAGIC (Su et al., 2022). This is because MAGIC fine-tuned the language model on the text corpus of MS-COCO captions. Compared to ZeroCap (Tewel et al., 2022b) and EPT (Tewel et al., 2022a), our CIDEr score improves 164.1% and 101.1%.
Table 8. Image captioning results on MS-COCO.
@@
ZeroCap (Tewel et al., 2022b) Zero-shot 2.9 12.0 - 13.1
MAGIC (Su et al., 2022) 12.9 17.4 - 49.3
EPT(Tewel et al., 2022a) 2.2 12.7 - 17.2
Ours 3.3 17.4 25.8 34.6
4.5. Qualitative Results

### v2

@@ 3.3 Inference
Optimization occurs throughout the sentence generation, enhancing the correspondence between the generated sentence and video through the iterative application. More specifically, the optimization is conducted at each step in the auto-regression process. When generating the stop token or reaching the maximum sentence generation length, we stop the process and initiate the generation of a new sentence while retaining the optimized soft prompt.
Figure 3: Our sentence-cleaning post-process strategy extract key elements, and use them to reconstruct the clean caption with the pre-trained CommonGen.
@@
After all $M$ iterations are complete, we obtain $M$ generated captions. We calculate the similarity between these $M$ generated captions and the video as follows:
$s^{*}=\arg\max_{s_{m}\in\{s_{1},s_{2},\ldots,s_{M}\}}\left(\frac{1}{T}\sum_{F\in\mathcal{F}}\operatorname{CLIP}(F,s_{m})\right).$ (10)
The caption $s^{*}$ with the highest similarity is selected as the reliable caption. However, due to the limitations of zero-shot learning, the selected caption still contains some noise and lacks fluency. We design a simple but effective sentence-cleaning strategy, which consists of two steps: (1) extracting essential elements like verbs and nouns from the selected sentence, and (2) employing a pre-trained keyword-sentence model CommonGen [34] to reconstruct these elements into a well-organized sentence. CommonGen is a state-of-the-art model for constrained text generation, which takes keywords (such as “knife”, “chopping”, “onion” and “kitchen”) as input and generates a more fluent sentence “a man is chopping onions with a knife in a kitchen”, as shown in Fig. 3. In the end, the cleaned sentence is used as the final result.
4 Experiments
4.1 Datasets and Metrics
@@
MSR-VTT [35] is a large-scale video captioning dataset collected from a commercial video website. It contains 10,000 video clips covering 20 categories. Each clip has 20 English annotations. In the standard split, it includes 6,513 training videos, 497 validation videos, and 2,990 test videos.
@@
MSVD [36] is a widely-use video captioning dataset, which is collected from Youtube. It consists of 1,970 video clips. Each clip has roughly 40 ground-truth captions. In the standard split, it includes 1,200 training videos, 100 validation videos, and 670 test videos.
@@
VATEX [37] is a large-scale dataset that contains 34,991 videos with 10 English annotations. The standard split includes 25,910 training videos, 3,000 validation videos, and 6,000 test videos.
@@
Metrics. In our experiments, we adopt four commonly used metrics to evaluate caption quality, including BLEU@4 (B@4) [38], METEOR (M) [39], ROUGE-L (R) [40], and CIDEr (C) [41]. Among these metrics, CIDEr is the main metric.
4.2 Implementation Details
@@
We use the pre-trained XCLIP [^1] https://huggingface.co/microsoft/xclip-base-patch16-kinetics-600-16-frames as our video retriever. We uniformly sample 16 frames per video for video retrieval. The visual feature extraction and text-image similarity calculation are performed using the pre-trained CLIP ViT-L/14222https://github.com/openai/CLIP model. We first sample three frames every second for the given video. To avoid using duplicate frames, we further use the CLIP-based strategy to subsample the keyframes. We start with the first frame as an anchor. The next frame is included in the set $\mathcal{F}$ if its dot product with the anchor, calculated by the CLIP image encoder, exceeds a specific threshold $\lambda_{CLIP}$. The newly sampled frame becomes the new anchor and the sampling process repeats in the same manner. The $\lambda_{CLIP}$ is set to 0.9 in all experiments. We use the pre-trained GPT-2 medium [^3] https://huggingface.co/openai-community/gpt2-medium as our caption generator, and the pre-trained AngLE [^4] https://github.com/SeanLee97/AnglE to calculate the similarity between the generated caption and retrieval sentences. Last, we leverage the pre-trained CommonGen [^5] https://huggingface.co/mrm8488/t5-base-finetuned-common_gen to clean the generated sentences into fluent ones. The aforementioned pre-trained models are publicly and readily available.
@@
Empirically, $\lambda_{l}$, $\lambda_{v}$, $\lambda_{s}$ and $\lambda_{w}$ are set to 1.6, 1.0, 0.8, and 0.3 respectively. In our method, we set the number of learnable tokens to 5. We set the number of retrieval sentences to 15 and the frequency size to 5. We adopt the AdamW [42] optimizer with a learning rate of 1e-4 and a weight decay weight of 0.3 to optimize our model. Throughout the experiments, we employ 16 generation iterations per video. To prevent sentences from being overly long and repetitive, we limit the number of tokens generated per sentence to 15. We implement our framework with PyTorch and all experiments are conducted on 1 $\times$ RTX-A6000 GPU.
@@
Table 1: Results on MSVD across various models and supervision levels.
Methods Setting Metrics
B@4 M R C
Results on MSVD test set
VNS-GRU [4] Supervised 66.5 42.1 79.7 121.5
SemSynAN [5] 64.4 41.9 79.5 111.5
HMN [43] 59.2 37.7 75.1 104.0
ZeroCap [13] Zero-shot 2.9 16.3 35.4 9.6
MAGIC [15] 6.6 16.1 40.1 14.0
EPT[14] 3.0 17.8 31.4 17.4
Video-LLaMA[44] 4.0 15.1 21.6 1.3
RETTA(Ours) 23.3 28.5 56.4 49.8
Table 2: Results on MSRVTT across various models and supervision levels.
@@
VNS-GRU [4] Supervised 45.3 29.9 63.4 53.0
SemSynAN [5] 46.4 30.4 64.7 51.9
HMN [43] 43.5 29.0 62.7 51.5
SGN [45] 40.8 28.3 60.8 49.5
SAAT [46] 40.5 28.2 60.9 49.1
ZeroCap [13] Zero-shot 2.3 12.9 30.4 5.8
MAGIC [15] 5.5 13.3 35.4 7.4
EPT[14] 3.0 14.6 27.7 11.3
Video-LLaMA[44] 4.9 16.8 25.3 2.3
ZS-CapCLIP[31] 4.0 15.0 31.0 5.0
MultiCapCLIP[18] 13.3 19.5 43.3 15.5
DeCap-BookCorpus [17] 6.0 12.7 - 12.3
DeCap-CC3M [17] 6.2 14.9 - 15.0
DeCap-COCO [17] 14.7 20.4 - 18.6
RETTA(Ours) 14.0 19.3 42.2 24.3
Table 3: Results on VATEX across various models and supervision levels.
@@
VaTeX [37] Supervised 28.4 21.7 - 45.1
HMN [43] 33.2 22.7 49.3 51.6
DeCap-BookCorpus [17] Zero-shot 4.1 9.9 - 11.8
DeCap-CC3M [17] 7.3 12.6 - 18.4
DeCap-COCO [17] 13.1 15.3 - 18.7
Video-LLaMA[44] 4.3 16.3 21.8 3.8
RETTA(Ours) 11.4 16.3 32.6 23.8
4.3 Performance Comparison
@@
We compare our method with several state-of-the-art models under different settings on three public benchmark datasets. The results are shown in Table 1, 2 and 3. We note that the performance of supervised algorithms is significantly higher than zero-shot methods, but they require expensive annotation costs. Compared to the other zero-shot methods, our method achieves the best performance in terms of the main metric CIDEr on all the three datasets. As shown in Table 1, our method achieves a CIDEr score of 49.8 on the MSVD benchmark, leading by about an absolute 32.4 over the previous best result. On the more challenging benchmark MSR-VTT, our method also boosts the CIDEr score by about 30.6%. On the VATEX dataset, our method achieves a 23.8 CIDEr Score, gaining 27.3% improvement in terms of the main metric CIDEr on VATEX.
4.4 Ablation Study
In this section, we perform a set of ablation studies on VATEX as it is the largest and most challenging.
Table 4: Performance with different our proposed component. Experiments are conducted on the VATEX dataset.
# $\mathcal{L}_{retrieval}^{S}$ $\mathcal{L}_{retrieval}^{W}$ $\mathcal{S}_{clean}$ B@4 M R C
1 1.4 9.8 19.4 8.9
2 2.5 13.3 22.6 20.8
3 2.6 12.2 22.1 14.9
4 2.4 13.5 22.9 21.4
5 11.4 16.3 32.6 23.8
Effect of our proposed component. To fully analyze the effect of each proposed component, we conduct experiments on the challenge VATEX. As shown in Table 4, we observe that each component assists the baseline to get better performance on all metrics. For example, Row 2 shows that using only sentence-granularity retrieval loss leads to a CIDEr score increase of $133.7\%$ compared to the baseline in Row 1. In contrast, employing word-granularity retrieval loss results in a CIDEr enhancement of $67.4\%$ and a BELU@4 score improvement of $85.7\%$ (Row 3 v.s. Row 1). The second-best performance is achieved when both types of losses are used together (Row 4 v.s. Row 1). In the comparison, the results with $\mathcal{S}_{clean}$ post-processing reach a further improvement of $11.2\%$ in the term of CIDEr (Row5 v.s.  Row4).
Effect of different retrieval corpora. In Table 5, we compare the performance of our method under different retrieval corpora. On the MSR-VTT benchmark, the CIDEr score increases around $9.9\%$ by leveraging the TestSet as the retrieval corpus instead of the TrainSet (Row 1 v.s. Row 2). On MSVD and VATEX datasets, employing TestSet improves around $21.9\%$ (Row 3 v.s. Row 4) and $8.4\%$ (Row 5 v.s. Row 6) compared to TrainSet. This demonstrates that different knowledge corpora can significantly impact performance, especially on relatively small datasets (e.g., MSVD). In this paper, unless otherwise specified, we use TrainSet as our default text corpus.
Effect of video retriever. In order to study the impact of video retrievers on our method, we use different configurations of XCLIP as our retriever. The results are present in Table 6. We observe that when XCLIP utilizes a ViT [47] with a size of B/16 as the backbone, it achieves an approximate improvement of 2.2% in the CIDEr score compared to using B/32 (Row 2 v.s. Row 1). The Kinetics-600 [25] dataset is the extended version of Kinetics-400 [48], containing more training data. Therefore, XCLIP pre-trained on Kinetics-600 brings a further improvement of 1.7% compared to XCLIP pre-trained on Kinetics-400 (Row 3 v.s. Row 2). Row 4 achieves the best performance on all metrics. It is trained with 16 frames per video and outperforms Row 3 using 8 frames. Therefore, in the other experiments conducted in this paper, we use the XCLIP in Row 4 as our default retriever.
@@ 4.4 Ablation Study
Effect of the number of retrieved sentences. We study the effect of using different numbers of retrieved sentences. Figure 4 shows the difference in video captioning performance as $K$ varies from 5 to 20. We find that retrieving 15 sentences offers better CIDEr scores and BLEU@4 scores compared to other numbers. It can be concluded that too many or too few retrieval sentences can degrade the performance. This is because too many retrieved sentences may introduce too much noise while too few retrieved sentences cannot provide sufficient information.
@@
Table 5: Effect of retrieval corpora. The experiments are conducted on test split of MSVD, MSR-VTT and VATEX datasets.
@@
MSR-VTT 1 TrainSet 14.0 19.3 42.2 24.3
2 TestSet 15.3 20.2 43.1 26.7
MSVD 3 TrainSet 23.3 28.5 56.4 49.8
4 TestSet 29.4 31.8 60.0 60.7
VATEX 5 TrainSet 11.4 16.3 32.6 23.8
6 TestSet 12.3 17.1 33.2 25.8
Table 6: Performance with different configuration XCLIP. The experiments are conducted on the VATEX dataset.
@@
1 VIT-B/32 8 Kinetics-400 10.5 15.6 31.9 22.3
2 VIT-B/16 8 Kinetics-400 10.9 15.8 32.2 22.8
3 VIT-B/16 8 Kinetics-600 11.2 16.1 32.3 23.2
4 VIT-B/16 16 Kinetics-600 11.4 16.3 32.6 23.8
@@
(a) CIDEr under different numbers of retrieved sentences
(b) BLEU@4 under different numbers of retrieved sentences
Figure 4: Performance (CIDEr and BLEU@4 score) comparison under different numbers of retrieved sentences on the VATEX dataset.
@@
Figure 5: Performance (CIDEr and BLEU@4 score) comparison under different numbers of soft tokens on the VATEX dataset.
Effect of the frequency size of the word-granularity loss. In Fig. 6, we conduct experiments on VATEX to explore the impact of the frequency size $L$ of word-granularity loss. From the figure, we observe that the CIDEr score first rises before reaching the moderate number (i.e., $L=5$), and then starts to decrease slightly. The main reason is that when $L$ is large, some noisy words are introduced. On the contrary, if $L$ is small, some essential words may be neglected. In Fig. 6(b), a similar trend is observed in the BLEU metric as well. Therefore, we select an appropriate number (i.e., $L=5$) as the frequency size.
@@
Figure 6: Performance (CIDEr and BLEU@4 score) comparison under different frequency sizes on the VATEX dataset.
Effect of the number of candidate words As mentioned in Sec. 3.2, to enhance efficiency, we compute the loss using the top 100 candidate words based on the original probability distribution obtained by the language model. In Figure 7(a) and Figure 7(b), we conduct experiments on VATEX to study the impact of the number of candidate words. From Figure 7(a), we find that using the top 250 candidate tokens achieves the highest CIDEr score compared to other numbers, with the top 100 candidate tokens achieving the second-best CIDEr score. In contrast, in Figure 7(b), we observe that the average time consumed per video increases roughly linearly with the number of candidate tokens. Compared to using 100 tokens, using 250 tokens takes about twice as long. To balance performance and time consumption, we set the number of candidate words to 100.
(a) CIDEr under different numbers of candidate words
(b) Average runtime under different numbers of candidate words
Figure 7: CIDEr and Average runtime per video under different numbers of candidate words on the VATEX dataset.
Effect of hard prompt set In our method, we randomly sample a phrase from a predefined set as the hard prompt. The set we designed includes four phrases: “Video showing”, “Video of”, “Video shows” and “Video describes”. In Table 7, we compare the performance of our method under different sizes of hard prompt sets. Row 1 displays the worst CIDEr score in the absence of any hard prompts. Using “Video showing” as a hard prompt consistently results in a CIDEr improvement of about 1.7% (Row 2 v.s. Row 1). The inclusion of “Video showing” and “Video of” in the hard prompt set leads to a CIDEr increase of around 0.8% (Row 3 v.s. Row 2). With the addition of “Video shows”, the CIDEr score in Row 4 leads by an absolute 0.3 over Row 3. Therefore, in our other experiments, the phrases covered in Row 4 are set as our default hard prompt set.
Table 7: Performance with different sizes of hard prompt sets. Experiments are conducted on VATEX.
# Hard Prompt Metrics
Video Video Video B@4 M R C
showing of shows
1 10.7 16.0 32.0 22.9
2 10.9 16.1 32.3 23.3
3 11.2 16.4 32.8 23.5
4 11.4 16.3 32.6 23.8
Zero-shot image captioning. Our method can be adapted for zero-shot image captioning by swapping XCLIP with CLIP. In Table 8, we compare the performance among several zero-shot image captions methods on the MS-COCO [49] dataset. We achieve the second-best results on all metrics, only falling behind MAGIC [15]. This is because MAGIC fine-tuned the language model on the text corpus of MS-COCO captions. Compared to ZeroCap [13] and EPT [14], our CIDEr score improves by an absolute 31.9 and 27.8, respectively.
Table 8: Image captioning results on MS-COCO.
@@
ZeroCap [13] Zero-shot 2.9 12.0 - 13.1
MAGIC [15] 12.9 17.4 - 49.3
EPT[14] 2.2 12.7 - 17.2
RETTA(Ours) 11.8 19.7 35.9 45.0
4.5 Qualitative Results